If the supply and demand functions are given by 20e0.4Q and P = 100e-0.2Q respectively, find the equilibrium price and quantity, and calculate the consumer’s and producer’s surplus



1
Expert's answer
2021-07-30T10:04:01-0400

Equating SUPPLY=DEMAND

Supply

"0.4Q=ln\frac{P}{20}"

"Q=\frac {ln\frac {P}{20}}{0.4}" …supply

"-0.2Q=ln\frac{P}{100}"

"Q=\frac{-ln\frac{P}{100}}{0.2}" …demand

"\frac {lnP-ln20}{0.4}=\frac {-lnP+ln100}{0.2}"

"P=96.15"

Substituting P in the equation for supply:

we get "Q=4."

Equilibrium price=96.15

Equilibrium quantity= 4